# MEGANet-W: A Wavelet-Driven Edge-Guided Attention Framework for Weak Boundary Polyp Detection

### v2

@@ Abstract
Colorectal polyp segmentation is critical for early detection of colorectal cancer, yet weak and low contrast boundaries significantly limit automated accuracy. Existing deep models either blur fine edge details or rely on handcrafted filters that perform poorly under variable imaging conditions. We propose MEGANet-W, a Wavelet Driven Edge Guided Attention Network that injects directional, parameter free Haar wavelet edge maps into each decoder stage to recalibrate semantic features. Our two main contributions are: (1) a two-level Haar wavelet head for multi orientation edge extraction; and (2) Wavelet Edge Guided Attention (W-EGA) modules that fuse wavelet cues with boundary and input branches. On five public polyp datasets, MEGANet-W consistently outperforms existing methods, improving mIoU by up to 2.3% and mDice by 1.2%, while introducing no additional learnable parameters.
1 Introduction
@@
U-Net along with variants, which are based on convolutional neural networks, have facilitated the rapid development of medical image segmentation tasks [29]. Subsequent studies use Transformer as encoder to strengthen the network’s global feature representation [32, 37]. Attention mechanism like SE-Net [9] and CBAM citeWoo2018 improve feature selectivity but offen blur boundaries due to spatial down-sampling, especially in low-contrast regions.
@@ 2 Related Work
Meanwhile, frequency-aware attention has gained traction for its ability to separate edge and context features. FAENet [25] apply discrete wavelet transform (DWT) to decompose features into high- and low-frequency bands, and WLAM [7] combines wavelet-derived low- and high- frequency features within a linear attention framework. Although promising, these designs are often limited to encoder-side processing or classification tasks, and few apply frequency-domain signals directly within decoder attention. Most closely related, WBANet [29] utilizes wavelet-based attention into a boundary-aware segmentation framework, yet lacks explicit edge-map injection into attention queries. Our work bridges this gap by feeding wavelet-derived edge features into decoder attention, enabling the model to adaptively focus on high-frequency boundary information during reconstruction, combining the strengths of spatial attention with the domain-adaptivity of wavelets.
3 Proposed Methodology
@@
3.1 Network Architecture
Our proposed network retains the encoder–decoder architecture of the [2] while integrating edge information at each decoding stage. Given an RGB image , a ResNet‐34 backbone extracts five spatial‐scale feature maps at strides , respectively. Starting from , a symmetric decoder with five upsampling blocks restores the input resolution. At each decoding stage , except the coarsest, we insert W-EGA module (see Fig. 2), replacing the Laplacian‐based EGA from [2]. Each decoder stage is projected to a side output via a standard ConvBlock (see Sec. 3.3), followed by a convolution. The side logits are bilinearly upsampled to the input resolution and supervised jointly.
3.2 Wavelet Enhanced Edge Head
@@
3.3 Wavelet Edge-Guided Attention
@@
Finally, a CBAM refresher recalibrates both channel and spatial responses, with a residual link preserving the original semantics:
@@ 3.4 Objective Function
4 Experiments
4.1 Datasets
@@ 4.3 Evaluation Metrics
5 Results
We quantitatively compare our MEGANet-W against seven state-of-the-art (SOTA) methods from the original MEGANet baseline paper: U-Net [21], U-Net++ [40], SFA [4], PraNet [6], MSNet [38], and PEFNet [19], as detailed in Sec.5.1. The results for MEGANet [2] were obtained by rerunning the model locally, while the results for all other methods are reported directly from the MEGANet paper. We also qualitatively compare our MEGANet-W with the original MEGANet baseline in Sec.5.2.
5.1 Quantitative Results
We evaluate MEGANet-W against a range of baselines. As shown in Table 1, our method consistently achieves the best or second-best performance across all datasets in terms of mIoU, mDice, and MAE. Notably, MEGANet-W outperforms all competing methods on CVC-300 with an mIoU of 82.8% and mDice of 89.5%, while also delivering SOTA results on Kvasir-SEG (87.0% mIoU, 92.0% mDice) and ETIS (69.6% mIoU, 77.5% mDice), confirming its robustness on both standard and challenging datasets. Across all datasets, MEGANet-W delivers the best MAE values, indicating improved pixel-level precision and better boundary localization. These results confirm the effectiveness of incorporating wavelet-based edge guidance within the decoder attention mechanism.
5.2 Qualitative Results
@@
6 Conclusion and Future Works
In this work, we presented MEGANet-W for weak-boundary polyp segmentation. Building upon the strong foundation of the original MEGANet [2] architecture, we integrate a parameter-free, two-level Haar wavelet head into each decoder stage and W-EGA module to fuse multi-orientation wavelet edges with reverse-attention and input-edge branches. MEGANet-W achieves more precise boundary delineation without increasing model complexity and consistently outperforms SOTA methods across five public polyp datasets in terms of mIoU, mDice, and MAE. Looking forward, futher work can focus on extending our work to video colonoscopy or volumetric data, paving the way for real-time tracking and 3D segmentation. Additionally, validating MEGANet-W on diverse, multi-center clinical datasets and enhancing its interpretability through visualization of wavelet attention responses will be essential steps toward clinical deployment and trust.
@@ Acknowledgements
We appreciate Xin Ying, Zhe Enn, and Zhe Yang for providing the computational units used in this work. We also thank Leon Loo and Zheng Yeh for their support.

### v3

@@ Abstract
Colorectal polyp segmentation is critical for early detection of colorectal cancer, yet weak and low contrast boundaries significantly limit automated accuracy. Existing deep models either blur fine edge details or rely on handcrafted filters that perform poorly under variable imaging conditions. We propose MEGANet-W, a Wavelet Driven Edge Guided Attention Network that injects directional, parameter free Haar wavelet edge maps into each decoder stage to recalibrate semantic features. The key novelties of MEGANet-W include a two-level Haar wavelet head for multi orientation edge extraction; and Wavelet Edge Guided Attention (W-EGA) modules that fuse wavelet cues with boundary and input branches. On five public polyp datasets, MEGANet-W consistently outperforms existing methods, improving mIoU by up to 2.3% and mDice by 1.2%, while introducing no additional learnable parameters. This approach improves reliability in difficult cases and offers a robust solution for medical image segmentation tasks requiring precise boundary detection.
Index Terms:
Colorectal Polyp Segmentation, Haar Wavelet Transform, W-EGA, Parameter-Free Attention, U-Net.I Introduction
@@
U-Net along with variants, which are based on convolutional neural networks, have facilitated the rapid development of medical image segmentation tasks [1]. Subsequent studies use Transformer as encoder to strengthen the network’s global feature representation [8, 9]. Attention mechanism like SE-Net [10] and CBAM [11] improve feature selectivity but offen blur boundaries due to spatial down-sampling, especially in low-contrast regions.
@@ II Related Work
Meanwhile, frequency-aware attention has gained traction for its ability to separate edge and context features. FAENet [12] apply DWT to decompose features into high- and low-frequency bands, and WLAM [33] combines wavelet-derived low- and high- frequency features within a linear attention framework. Although promising, these designs are often limited to encoder-side processing or classification tasks, and few apply frequency-domain signals directly within decoder attention. Most closely related, WBANet [1] utilizes wavelet-based attention into a boundary-aware segmentation framework, yet lacks explicit edge-map injection into attention queries. Our work bridges this gap by feeding wavelet-derived edge features into decoder attention, enabling the model to adaptively focus on high-frequency boundary information during reconstruction, combining the strengths of spatial attention with the domain-adaptivity of wavelets.
III Proposed Methodology
@@
III-A Network Architecture
Our proposed network retains the encoder–decoder architecture of the [5] while integrating edge information at each decoding stage. Given an RGB image , a ResNet‐34 backbone extracts five spatial‐scale feature maps at strides , respectively. Starting from , a symmetric decoder with five upsampling blocks restores the input resolution. At each decoding stage , except the coarsest, we insert W-EGA module (see Fig. 2), replacing the Laplacian‐based EGA from [5]. Each decoder stage is projected to a side output via a standard ConvBlock (see Sec. III-C), followed by a convolution. The side logits are bilinearly upsampled to the input resolution and supervised jointly.
III-B Wavelet Enhanced Edge Head
@@
III-C Wavelet Edge-Guided Attention
@@
Finally, a CBAM refresher consists of sequential channel and spatial attention modules, which focus on the channel and spatial dimensions, respectively. This mechanism recalibrates feature responses, while a residual connection preserves the original semantics, similar to the attention strategy employed in [5].:
@@ III-D Objective Function
* Reproduced using official code
IV Experiments
IV-A Datasets
@@ IV-C Evaluation Metrics
V Results
We quantitatively compare our MEGANet-W against seven state-of-the-art (SOTA) methods from the original MEGANet baseline paper: U-Net [34], U-Net++ [35], SFA [14], PraNet [15], MSNet [28], and PEFNet [31], as detailed in Sec.V-A. The results for MEGANet [5] were obtained by rerunning the model locally, while the results for all other methods are reported directly from the MEGANet paper. We also qualitatively compare our MEGANet-W with the original MEGANet baseline in Sec.V-B.
V-A Quantitative Results
We evaluate MEGANet-W against a range of baselines. As shown in Table I, our method consistently achieves the best or second-best performance across all datasets in terms of mIoU, mDice, and MAE. Notably, MEGANet-W outperforms all competing methods on CVC-300 with an mIoU of 82.8% and mDice of 89.5%, while also delivering SOTA results on Kvasir-SEG (87.0% mIoU, 92.0% mDice) and ETIS (69.6% mIoU, 77.5% mDice), confirming its robustness on both standard and challenging datasets. Across all datasets, MEGANet-W delivers the best MAE values, indicating improved pixel-level precision and better boundary localization. These results confirm the effectiveness of incorporating wavelet-based edge guidance within the decoder attention mechanism.
V-B Qualitative Results
@@
VI Conclusion and Future Works
In this work, we presented MEGANet-W for weak-boundary polyp segmentation. Building upon the strong foundation of the original MEGANet architecture, we integrate a parameter-free, two-level Haar wavelet head into each decoder stage and W-EGA module to fuse multi-orientation wavelet edges with reverse-attention and input-edge branches. MEGANet-W achieves more precise boundary delineation without increasing model complexity and consistently outperforms SOTA methods across five public polyp datasets in terms of mIoU, mDice, and MAE. Looking forward, futher work can focus on extending our work to video colonoscopy or volumetric data, paving the way for real-time tracking and 3D segmentation. Additionally, validating MEGANet-W on diverse, multi-center clinical datasets and enhancing its interpretability through visualization of wavelet attention responses will be essential steps toward clinical deployment and trust.
@@ Acknowledgements
We appreciate Xin Ying, Zhe Enn, and Zhe Yang for providing the computational units used in this work. We also thank Leon Loo and Zheng Yeh for their support. This research work is supported by Xiamen University Malaysia Research Fund (Grant No: XMUMRF/2024-C14/IECE/0055).